OpenStudy (anonymous):

what is (1/4)^3 and 3(1/4)

OpenStudy (anonymous):

\[(1/4)^3 = 1^3 / 4^3\ = 1/64\] and 3(1/4)=(3x1)/4 = 3/4

OpenStudy (anonymous):

tnx bfsgd and pls what about (√2)^3?

OpenStudy (anonymous):

\[(\sqrt{2})^3 = \sqrt{2} \times \sqrt{2} \times \sqrt{2} = 2 \sqrt{2} \]

OpenStudy (anonymous):

and 3(√2)?

OpenStudy (anonymous):

Like the previous example that can't be simplified any further. It's just 3 times the square root of 2.